100cm3 of 0.20 mol/dm3 K2SO4 are added to 100cm3 of water and mixed thoroughly. What is the molarity of K+ ions in the resulting solution.


1
Expert's answer
2021-01-25T04:35:36-0500

Initial Volume(V1) = 100 cm3 = 100/1000 = 0.10 dm3

Final Volume (V2) = 100 +100 = 200 cm3 = 200/1000 = 0.20 dm3

initial concentration of K2SO4 (M1)= 0.20 mol/dm3

Now, Using dilution law,

M1V1 = M2V2

0.10 × 0.10 = M2× 0.20

M= 0.10 × 0.10 /0.20

= 0.050 mol/dm3

Now, concentration of K+ = 2×0.050 = 0.10 mol/dm3